Show that the function f defined by

F(x)=x^3+4x^2+x-6

has a real root in the interval [0,2]


1
Expert's answer
2021-05-10T18:23:01-0400

Solution.


"f(x)=x^3+4x^2+x-6"

Domain(f)="(-\infty,\infty)."

"f(0)=-6<0."

"f(2)=8+16+2-6=20>0."

There are must be at least one real root between 0 and 2.




Check point "x=1, 1 \in (0,2)."

"f(1)=1+4+1-6=0."

So, "x=1" is real root of function f(x) on interval (0,2).